OpenStudy (anonymous):

$61.00 with a 7% sales tax

OpenStudy (anonymous):

61.00 + (7% of 61) = 61.00 + (0.07*61) = 61.00 + 4.27 = $65.27

OpenStudy (ttop0816):

61.00 + (7% of 61) = 61.00 + (0.07*61) = 61.00 + 4.27 = $65.27

OpenStudy (anonymous):

@ttop0816 um

OpenStudy (anonymous):

thx